OpenStudy (anonymous):

←→ ←→ AB and BC form a right angle at point B. If A = (-3, -1) and ←→ B = (4, 4), what is the equation of BC? (Please Help with FAN and Medal!)

OpenStudy (anonymous):

A.) x + 3y = 16 B.) 2x + y = 12 C.) -7x − 5y = -48 D.) 7x − 5y = 48